Circular No. 6787

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 971214 J. Heise and J. in ’t Zand, Space Research Organization, Utrecht; G.

Spoliti, V. Torroni, and D. Ricci, BeppoSAX, Rome; L. Piro, E. Costa, and M. Feroci, Instituto di Astrophysica Spaziale, Rome; and F. Frontera, Instituto di Tecnologie e Studio delle Radiazoni Extraterrestri, Bologna, report: “The BeppoSAX Gamma Ray Burst Monitor was triggered on Dec. 14.97272 UT. A quick-look analysis shows a structured γ-ray burst lasting ∼ 25 s with a peak flux of 650 counts/s. The γ-ray burst was also detected by the Wide Field Camera WFC1 with a peak flux of ∼ 1 Crab. The position from a quick-look analysis is α = 11^{h}56^{m}30^{s}, δ = +65^{o}12′.0 (equinox 2000.0), with an error radius of 3′.9 (99-percent confidence). A BeppoSAX Narrow Field follow-up observation is in progress. Observations at all wavelengths are urged.”

PSR J0633+1746 A. Shearer, A. Golden, and R. M. Redfern, National University of Ire-

land, Galway; and G. M. Beskin, Special Astrophysical Observatory, write: “We observed PSR J0633+1746 (Geminga) on 1996 Jan. 12–14 with the Russian 6-m telescope (+ TRIFFID camera + MAMA detector) in bands B (29 307 s) and V (6829 s). We found optical pulsations in B with magnitude 26.0 ± 0.4 and a 3σ upper limit for pulsations in V of ∼ 25.5. Enhanced sky brightness limited a detection in V. The significance of the B result using the phase-independent Z statistic is 99.96 percent, using the ephemeris by Mattox et al. (1996, A.Ap. Suppl. 120, C77). The corresponding lightcurve shows two peaks in phase with γ-ray and hard-x-ray data, testifying to a dominant non-thermal component in B. Deeper timeresolved exposures in B, V, R, and I should confirm pulsed activity in these other wavebands, and provide important data as regards understanding the thermal and nonthermal components of Geminga’s emission.”

COMET C/1997 J2 (MEUNIER-DUPOUY) Visual m_{1} estimates: Nov. 17.75 UT, 11.3 (R. J. Bouma, Groningen,

The Netherlands, 0.25-m refl.); 27.11, 10.8 (C. E. Spratt, Victoria, BC, 0.20m refl.); Dec. 2.98, 11.7 (J. Bortle, Stormville, NY, 0.41-m refl.); 14.72, 9.9 (M. Lehký, Hradec Králové, Czech Rep., 25×100 binoculars).

1997 December 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green